OpenStudy (prestonsetzer1):

If (x + 1)(x - 3) = 5, then which of the following statements is true? x + 1 = 0 or x - 3 = 0 x + 1 = 5 or x - 3 = 5 x - 4 = 0 or x + 2 = 0

alones (alones):

Simplify both sides of the equation. \[x^2−2x−3=5\] and subtract 5 from both sides \[x^2−2x−3−5=5−5\] \[x^2−2x−8=0\] Factor left side of equation. \[(x+2)(x−4)=0 \] Set factors equal to 0 \[x+2=0 or x−4=0 \] which would bring us to\( x=−2\) or \(\ x=4 \)

OpenStudy (awhite):

This isn't FIOL right?

alones (alones):

So basically the answer should be \(x - 4 = 0\)or \(x + 2 = 0 \)

OpenStudy (mathstudent55):

There was FOIL used in the first step.

OpenStudy (mathstudent55):

All 3 choices are incorrect. Are you sure you copied the second choices correctly?

OpenStudy (mathstudent55):

@calculusxy That is not correct.

OpenStudy (calculusxy):

How are all three incorrect?

OpenStudy (mathstudent55):

Sorry, you are correct. One choice is correct. It's the one that gives the correct solutions that @AloneS found.

alones (alones):

Wait so i'm wrong?

OpenStudy (mathstudent55):

One of the three choices of the original problem that was posted is correct.

OpenStudy (mathstudent55):

@AloneS is correct. One of the choices is @AloneS 's solution.

OpenStudy (mathstudent55):

What is incorrect is a statement that @calculusxy made above.

OpenStudy (calculusxy):

How can you confirm that the option works?

OpenStudy (mathstudent55):

If (x + 1)(x - 2) = 0, then one of the factors must equal zero. That is how we solve the equation. If (x + 1)(x - 3) = 5, we cannot set the factors equal to 5 and solve. That simply does not work.

OpenStudy (calculusxy):

I see

OpenStudy (mathstudent55):

Look at the solution we have: x = 4 or x = -2 One of the choices is the same as that solution.

OpenStudy (calculusxy):